# PRISM: A Framework Harnessing Unsupervised Visual Representations and Textual Prompts for Explainable MACE Survival Prediction from Cardiac Cine MRI

### v1

@@ 2.2 Visual signatures of MACE in cardiac motility
(a) Attention rollout of image tokens in Stage II across cardiac cycle phases.
(b) Spatiotemporal KDEs of predicted heatmap distributions in the SAX view.
Figure 5: Model attention and heatmap analysis across cardiac cycle phases and MACE risk. (a): Attention rollout of image tokens under stage II with horizontal axis representing typical cardiac cycle phases (early diastole, late diastole, ventricular ejection, isovolumetric relaxation) and vertical axis indicating increasing MACE risk levels. Text spans highlighted in light blue showing strong concordance between model attention and radiological diagnosis. (b), Spatiotemporally separated kernel density estimates of model-predicted heatmap distributions across three risk levels, six myocardial segments (A, I, AS, IS, AL, IL), and four cardiac cycle periods (early diastole, late diastole, ventricular ejection, isovolumetric relaxation), in the SAX view.
@@ 2.3 BiPromptSurv-driven EHR analysis
Intuitively, comprehensive EHR integration yields the most accurate MACE risk stratification. However, our BiPromptSurv framework empirically identifies cohort-preferred EHR feature groups, enabling prompt-guided secondary survival analysis. By isolating salient feature subsets, this approach improves risk stratification consistency under controlled conditions, underscoring the prognostic value of selective feature integration. We introduced a set of syntactically diverse prompts absent from the model’s pre-training corpus. The prompts were synthesized via DeepSeek-R1 70B [41], generating a prompt distribution, which extrapolated novel yet semantically aligned instructions from a small set of handcrafted exemplars. A lightweight routing module assigns these prompts to categorized EHR class weights.
@@
Among the generated prompts, a handful exhibited consistent improvements in risk stratification fidelity across target cohorts under the BiPromptSurv framework’s instruction for secondary survival analysis. Notably, as illustrated in Fig. 6a, prompting with statements that focus on a single EHR category, such as Design your metrics pipeline around <clinical>, and using only 13 clinical-related features leads to a 1.2% increase in the mean AUC for the RJCCM cohort, albeit at the cost of higher cross-seed variance. Similarly, applying prompts like Incorporate multimodal indicators from <clinical> and <physiological> information to enhance the precision of the machine learning model led to improved survival prediction in both GLCCM and RJCCM cohorts by leveraging clinical and physiological features, outperforming models trained on the full feature set.
(a) Performance gains with BiPromptSurv pipeline and comparison with all-EHR baseline.
(b) CoxPH-based cluster attribution of EHR categories for survival impact.
(c) SHAP attribution of PRISM highlighting top-ranked EHR and image features.
@@
Figure 6: Statistical analysis results based on PRISM Stage II and Stage III. (a): Examples of prompts exhibiting performance improvement under the BiPromptSurv pipeline along with result comparisons. Dot plots with error bars in light blue and light green across five seeds are displayed. Red arrows indicate the increase in mean corresponding metrics compared to the all EHR setting. (b): Cluster attribution analysis based on CoxPH, showing contributions of each EHR category reflecting their survival analysis impact within the cohort. (c): Fine-grained SHAP attribution analysis of PRISM. Top-10 ranked features are presented along with top-5 image features aggregated as Aggregated_Img_feature.
@@
Building on this foundation, CoxPH model was employed, which serves as the gold standard and as the survival head within our framework, to conduct cohort-specific attribution analysis on the selected feature groups. Aggregated SHAP values were derived to demonstrate the contributions of different types of EHR features. As illustrated in Fig. 6b, CoxPH attribution reveals that clinical and physiological features dominate the EHR contribution in RJCCM and GLCCM cohorts, accounting for approximately 82.4% and 69.9% of the total SHAP values, respectively. This concordance confirms BiPromptSurv’s targeted feature selection efficacy while underscoring the interpretability of prompt-guided imaging features.
@@
Furthermore, SHAP analysis was employed for fine-grained attribution, aiming to uncover how our model’s Stage II and III behaviors vary across different prompts and the respective contributions of specific categorized EHR features to survival analysis. Using PRISM, we identified and presented the top 10 most important features at the cohort level, among which is an Aggregated_Img_feature aggregated from the five core imaging features. Concordance was observed in GLCCM, where physiological indicators such as heart rate (HR), stroke volume (SV), and ejection fraction (EF), which are closely related to myocardial function, were included. This is consistent with the role of PRISM in extracting imaging features during Stage I. Meanwhile, clinical features such as hypertension, diabetes, and dyslipidemia also remained dominant. The same trend was observed in RJCCM, where physiological and clinical features accounted for 80% of the top 10. Additionally, the role of aggregated image features as the top-ranked primary contributors in the TJCCM cohort is prominently demonstrated in Fig. 3, further validating the effectiveness of the PRISM model in image analysis by achieving significant improvements over classical EHR models, SOTA methods, and unsupervised comparative approaches.
@@ 2.4 Individualized PRISM-based risk interpretation
In contrast, among the listed complications strongly associated with MACE, case B exhibited two conditions, including microcirculatory disorder and ventricular aneurysm. Compared to MACE=0 samples, its differential EHR feature ranking showed less pronounced divergence than in case A, as reflected in the smaller Z-score deviation from zero. This observation may partly explain why PRISM adopted a more conservative approach in predicting MACE for case B.
@@ 3 Discussion
Integrating the cross-cohort SHAP value analyses derived from our PRISM model, we identified that among routinely obtained clinical assessments in standard hospital settings, hypertension and diabetes emerge as prominent potential biomarkers with high priority for predicting MACE. Notably, these comorbidities consistently ranked within the top 10 salient features across three of our independent centers. This finding underscores the persistent and critical role of traditional EHR variables alongside the advanced, unsupervised, image-based feature extraction facilitated by PRISM in survival analyses.
@@
4 Methods
4.1 Study Population and Inclusion Criteria
A total of 1970 patients were retrospectively analyzed from four tertiary medical centers in this study. Eligibility required confirmation of first-onset ST-segment elevation myocardial infarction through characteristic ischemic symptoms accompanied by significant ST-segment elevation (>> 0.1 mV in limb leads or >> 0.2 mV in precordial leads) in at least two contiguous leads. All enrolled cases received primary percutaneous coronary intervention within 12 hours of symptom onset and subsequently underwent cardiovascular magnetic resonance examination within 7 days following reperfusion. Exclusion criteria comprised previous myocardial infarction, contraindications to cardiovascular magnetic resonance, suboptimal imaging quality, or established non-ischemic cardiomyopathy diagnoses.
4.2 Follow-up Protocol and Endpoint Assessment
Patient follow-up commenced upon discharge following the index event and continued through December 31, 2024. Outcome data were systematically obtained through structured telephone interviews, standardized questionnaires, and scheduled clinical evaluations conducted at three-month intervals.
The primary composite endpoint of MACE incorporated four components: (1) mortality from any cause, (2) recurrent myocardial infarction as per the Fourth Universal Definition of Myocardial Infarction (2018), (3) unscheduled revascularization procedures, and (4) hospital admissions due to heart failure.
Heart failure diagnoses required either: (a) expert clinical confirmation or (b) documentation of new-onset symptomatic fluid retention necessitating diuretic therapy in patients demonstrating either left ventricular ejection fraction below 50% or elevated natriuretic peptide levels. Revascularization procedures were classified as unplanned if they were neither part of the initial therapeutic strategy nor scheduled to occur within 60 days post-index procedure.
All events occurring prior to cardiovascular magnetic resonance imaging were excluded from final analysis. An independent central laboratory conducted blinded adjudication of all reported clinical endpoints to ensure standardized event classification.
4.3 Image preprocessing: ventricular motion focus
Cardiac cine MRI sequences underwent mask-free motion analysis to quantify ventricular dynamics without contrast agents. Full cardiac volumes $\mathcal{V}^{(total)}$ were processed through Equ. 1 using Farnebäck optical flow [47] to compute pixel-wise displacement vectors and capture inter-frame myocardial motion,
where $\mathbf{F}^{(t)}$ represents the computed motion vector field between consecutive frames $\mathcal{I}_{t}$ and $\mathcal{I}_{t+1}$, with $\Psi_{\mathrm{FB}}$ denoting the optical flow operator.
The global region-of-interest was centered on ventricular motion trajectories, with a standardized 96-pixel window centered on the cardiac chambers. This approach selectively enhances visualization of biventricular wall motion while eliminating confounding signals from extracardiac structures, enabling precise quantification of regional contractility abnormalities without manual segmentation.
4.4 Stage I: motion-aware multi-view distillation
The PRISM’s knowledge distillation framework gains insights from the DINOv2 [48] paradigm in natural image processing, adapting its self-supervised principles to multi-view cardiac imaging through asymmetrical processing pathways.
The teacher pathway integrates long-axis views comprising 2-chamber, 3-chamber and 4-chamber orientations, capturing comprehensive myocardial dynamics across multiple anatomical planes. Simultaneously, the student pathway analyzes spatially compressed short-axis sequences through dimensionality-reduced representations, learning condensed motion features from the cine SA data. Feature alignment between pathways is achieved via a distillation optimization expressed in Equ. 2,
where the teacher’s multi-view feature embeddings, specifically denoted as $\bar{p}^{(t)}$, guide the student’s SA-derived representations $p^{(s)}$ through Kullback-Leibler divergence [49] minimization. This mechanism transfers knowledge regarding cross-view myocardial motion patterns to the compressed SA pathway. The complementary contrastive component, identified as $\mathcal{L}_{\text{contrastive}}$ and instantiated via an InfoNCE-based formulation [50], coordinates phase-specific feature consistency within individual patients while simultaneously differentiating motion characteristics across distinct subjects. This dual pathway strategy yields spatiotemporally coherent motion tokens encoding prognostic signatures characteristic of ventricular dysfunction.
4.5 Stage II: EHR-attention directive
Given a medical prompt $\mathcal{T}\in\mathbb{V}^{L}$, where $\mathbb{V}$ denotes the clinical vocabulary and $L$ the sequence length, the routing module $\mathcal{R}$ generates a sparse group attention vector $\boldsymbol{\alpha}=[\alpha_{1},\dots,\alpha_{G}]$ through semantic parsing expressed in Equ. 3,
$\boldsymbol{\alpha}=\mathcal{R}(\text{Embed}(\mathcal{T})),$ (3)
where each $\alpha_{i}$ is approximately 0, 0.5, or 1.
The architecture then branches into three parallel processing streams to establish cross-modal alignment.
The Cross-Attention Fusion component computes query projections $\mathbf{Q}=\mathbf{P}\mathbf{W}_{Q}\in\mathbb{R}^{N_{P}\times d_{attn}}$ from BERT-encoded [51] prompt embeddings $\mathbf{P}\in\mathbb{R}^{N_{P}\times d_{P}}$, while spatial-temporal visual features $\mathbf{Z}^{(s)}\in\mathbb{R}^{N_{I}\times d_{I}}$ generate keys and values via linear projection $\mathbf{K}=\mathbf{V}=\mathbf{Z}^{(s)}\mathbf{W}_{K}$. The aligned image representation emerges through multi-head attention followed by dimension compression in Equ. 4,
where $\oplus$ denotes concatenation along the sequence dimension and $\mathbb{E}_{l}[\cdot]$ implements average pooling over the prompt token axis. $\mathbf{Q}_{l}$ is a prompt query constructed in an inference-intuitive manner based on given medical prompt.
For EHR feature processing, the Embedding Engine converts categorical EHR variables $\mathbf{e}\in\mathbb{Z}^{m}$ into group-specific representations. Each feature $e_{f}$ in group $g$ is embedded through dedicated layers $\phi_{f}(\cdot)$, then aggregated with routing weights $\alpha_{g}$ demonstrated in Equ. 5,
where $S_{g}$ indexes clinical feature subgroups. Finally, the Visual Anchor stream preserves original feature topology through detached projection of spatial-temporal features outlined in Equ. 6,
$\mathbf{Z}_{ref}=\text{LayerNorm}\left(\mathbf{W}_{r}\cdot\mathbb{E}[\mathbf{Z}^{(s)}]\right).$ (6)
4.6 Alignment Objectives
where $\mathcal{D}_{SW}$ denotes sliced Wasserstein distance. Then we obtain the composite objective in Equ. 9,
$\mathcal{L}_{total}=\mathcal{L}_{tri}+\beta\mathcal{L}_{pres}.$ (9)
4.7 Stage III: survival prediction
Clinical indicators $\mathbf{e}\in\mathbb{R}^{d_{m}}$ derived from tabulated EHR provide essential information for constructing a multimodal fusion module together with pre-trained image tokens $\mathbf{r}=\mathbb{E}(Q_{img})\in\mathbb{R}^{d_{c}}$. These features are concatenated to form the combined vector $\mathbf{x}_{\mathrm{fused}}=\mathrm{concat}[\mathbf{e},\mathbf{r}]\in\mathbb{R}^{d_{m}+d_{c}}$, integrating informative data from both clinical and imaging domains.
A standard CoxPH head is applied, where the coefficients $\beta_{k}$ serve to automatically weight the contributions of each fused feature. The hazard function is modeled in Equ. 10,
with $h_{0}(t)$ representing the baseline hazard function.
The training objective minimizes the negative log-partial likelihood given by Equ. 11,
where $\mathbf{\theta}=[\beta_{1},\ldots,\beta_{m}]^{\top}$ denotes the vector of regression parameters, $\delta_{i}$ indicates whether the event occurred for subject $i$, and $R(t_{i})$ corresponds to the risk set at survival time $t_{i}$.
4.8 Evaluation metrics
In this study, we utilized C-index [52] and AUC as primary evaluation metrics. The C-index assesses the discriminatory power of the model by measuring the proportion of all comparable subject pairs whose predicted risks are correctly ordered relative to their observed survival times. Formally, it is defined in Equ. 12,
where $\delta_{i}$ indicates the occurrence of the event of interest, and $r_{i}=\mathbf{\beta}^{\mathrm{T}}\mathrm{x}_{\mathrm{fused}}^{i}$ represents the risk score derived from the prognostic model. Complementarily, we incorporated AUC to quantify the model’s sensitivity and specificity in stratifying patients with respect to event status across varying decision thresholds.
4.9 Implementation details
The model was implemented in the PyTorch environment and trained using NVIDIA-H100 GPUs with CUDA 12.2. Specifically, Stages I and III were conducted on a single H100 GPU, whereas Stage II utilized a 4-GPU H100 setup. The GPU-to-CPU ratio was maintained at approximately 1 to 15. Optimization was performed using the Adam optimizer [53] with an initial learning rate of $5\times 10^{-5}$ and weight decay set to $1\times 10^{-5}$. A batch size of 16 was used throughout, and models were trained for 50 epochs with a StepLR learning rate scheduler. The 4D cardiac cine MRI inputs were uniformly resized to dimensions of $24\times 24\times 96\times 96$, indicating depth, temporal frames, height, and width. A total of 41 EHR features were incorporated, with 579, 672, 406, and 313 samples included from GLCCM, AZCCM, RJCCM, and TJCCM, respectively.
Standard data partitioning schemes were employed, with a training-to-validation ratio of 7:3 for self-supervised learning during Stages I and II, and a split of 6:2:2 for training, validation, and testing sets in the survival prediction stage (Stage III). In Stage II, each image sample was paired with 50 unique medical prompt variants. During Stage III, grid search combined with stratified sampling was used for hyperparameter tuning, and feature selection was performed using stepwise regression and Lasso regularization within the CoxPH. The IECV setting is implemented by fixing the same test set within each cohort. For internal validation, training and testing are performed on the individual cohort separately. For external validation, the model is trained on the combined training sets of the remaining cohorts and tested on the same test set as used in the internal validation.
@@ 4.9 Implementation details
The Institutional Review Boards (IRBs) of all participating centers approved this research. To protect patient privacy, informed consent was waived by the IRBs, and all data were anonymized prior to their use in model training, testing, and reader studies.
@@
H.S. designed the model and performed experiments. H.S. and S.R. developed and provided survival analysis code. H.S., S.R., and Y.G. processed the experimental data and documented control experiment details. J.X., X.C., T.Y., and L.W. provided cine imaging data, medical diagnoses, and electronic health records. L.W. and X.W. supervised the research direction. H.S. drafted the manuscript. J.X., L.W., and X.W. contributed to manuscript editing and revision.

### v2

@@ 3.2 Visual signatures of MACE in cardiac motility
Figure 6: Model attention and heatmap analysis across cardiac cycle phases and MACE risk. (a), Attention rollout of image tokens under stage II with horizontal axis representing typical cardiac cycle phases (early diastole, late diastole, ventricular ejection, isovolumetric relaxation) and vertical axis indicating increasing MACE risk levels. Text spans highlighted in light blue showing strong concordance between model attention and radiological diagnosis. (b), Spatiotemporally separated kernel density estimates of model-predicted heatmap distributions across three risk levels, six myocardial segments (A, I, AS, IS, AL, IL), and four cardiac cycle periods (early diastole, late diastole, ventricular ejection, isovolumetric relaxation), in the SAX view.
@@ 3.3 BiPromptSurv-driven EHR analysis
Intuitively, comprehensive EHR integration yields the most accurate MA-CE risk stratification. However, our BiPromptSurv framework empirically identifies cohort-preferred EHR feature groups, enabling prompt-guided secondary survival analysis. By isolating salient feature subsets, this approach improves risk stratification consistency under controlled conditions, underscoring the prognostic value of selective feature integration. We introduced a set of syntactically diverse prompts (Supplementary Fig. A.1) absent from the model’s pre-training corpus. The prompts were synthesized via DeepSeek-R1 70B [7], generating a prompt distribution, which extrapolated novel yet semantically aligned instructions from a small set of handcrafted exemplars. A lightweight routing module assigns these prompts to categorized EHR class weights. The routing module’s performance in recognizing different medical semantics of EHR categories is presented in the UpSet plot in Supplementary Fig. A.2.
@@
Within the BiPromptSurv framework, a subset of generated prompts consistently improved risk stratification fidelity when applied to secondary survival analysis. Notably, as illustrated in Fig. 7, prompting with statements that focus on a single EHR category, such as Design your metrics pipeline around <clinical>, leads to a 1.2% increase in the mean AUC for the RJCCM cohort, albeit at the cost of higher cross-seed variance. Similarly, applying prompt like Incorporate multimodal indicators from<clinical> and <physiological> information to enhance the prec-ision of the machine learning model led to improved survival prediction in both GLCCM and RJCCM cohorts by leveraging clinical and physiological features, outperforming models trained on the full feature set.
@@
Figure 7: Examples of prompts under the BiPromptSurv pipeline along with result comparisons. Dot plots with error bars in light blue and light green across five seeds are displayed. Red arrows indicate the increase in mean corresponding metrics compared to the all EHR settings.
Figure 8: Statistical analysis results based on PRISM Stage II and Stage III. (a), Cluster attribution analysis based on CoxPH, showing contributions of each EHR category reflecting their survival analysis impact within the cohort. (b), Fine-grained SHAP attribution analysis of PRISM. Top-10 ranked features are presented along with top-5 image features aggregated as Aggregated_Img_feature.
@@
Building on this foundation, CoxPH model was employed, which serves as the gold standard and as the survival head within our framework, to conduct cohort-specific attribution analysis on the selected feature groups. Aggregated SHAP values were derived to demonstrate the contributions of different types of EHR features. As illustrated in Fig. 8a, CoxPH attribution reveals that clinical and physiological features dominate the EHR contribution in RJCCM and GLCCM cohorts, accounting for approximately 82.4% and 69.9% of the total SHAP values, respectively. This concordance confirms BiPromptSurv’s targeted feature selection efficacy while underscoring the interpretability of prompt-guided imaging features.
@@
Furthermore, SHAP analysis was employed for fine-grained attribution in Fig. 8b, aiming to uncover how our model’s Stage II and III behaviors vary across different prompts and the respective contributions of specific categorized EHR features to survival analysis (the full EHR-level SHAP analysis is provided in Supplementary Fig. A.3). Using PRISM, we identified and presented the top-10 most important features at the cohort level, among which is an Aggregated_Img_feature aggregated from the five core imaging features. Concordance was observed in GLCCM, where physiological indicators such as heart rate (HR), stroke volume (SV), and ejection fraction (EF), which are closely related to myocardial function, were included. This is consistent with the role of PRISM in extracting imaging features during Stage I. Meanwhile, clinical features such as hypertension, diabetes, and dyslipidemia also remained dominant. The same trend was observed in RJCCM, where physiological and clinical features accounted for 80% of the top-10. Additionally, the role of aggregated image features as the top-ranked primary contributors in the TJCCM cohort is prominently demonstrated in Fig. 4, further validating the effectiveness of the PRISM model in image analysis by achieving significant improvements over classical EHR models, SOTA methods, and unsupervised comparative approaches.
3.4 Ablation Study
Table 3 and Table 4 present the ablation study results comparing PRISM with and without Stage II across all cohorts under IECV settings. PRISM without Stage I represents an ablation version in which SSL distillation is not performed, and the original UniFormer weights pre-trained on general domain are directly used for Stage II. PRISM without Stage II represents an ablation version of our proposed model in which the visual features from Stage I are directly fed into survival analysis.
Table 3: Ablation study results (C-index) comparing PRISM with and without Stage II across cohorts under IECV settings.
Model GLCCM AZCCM RJCCM TJCCM
Internal
[…]
External
[…]
Table 4: Ablation study results (AUC) comparing PRISM with and without Stage II across cohorts under IECV settings.
Model GLCCM AZCCM RJCCM TJCCM
Internal
[…]
External
[…]
To enable PRISM modeling in the out-domain setting, we harmonized the categorical structure of EHR features and examined the effect of applying or omitting strata stratification, allowing the CoxPH survival head to adaptively learn a set of basis functions from the training data. As shown in Table 5 and Table 6, we systematically evaluated different feature selection strategies (stepwise regression and Lasso regularization) combined with strata stratification across all four cohorts under external validation settings.
Table 5: Survival analysis results of the PRISM model using prompt defined in the main table, evaluated under different feature selection strategies and strata setting. Abbreviations: Str, strata; Ste, stepwise; L, lasso. Metrics: C-index.
Str Ste L GLCCM C-index AZCCM C-index RJCCM C-index TJCCM C-index
[…]
Table 6: Survival analysis results of the PRISM model using prompt defined in the main table, evaluated under different feature selection strategies and strata setting. Abbreviations: Str, strata; Ste, stepwise; L, lasso. Metrics: AUC.
Str Ste L GLCCM AUC AZCCM AUC RJCCM AUC TJCCM AUC
[…]
We observed that enabling strata stratification consistently improved the survival analysis performance of PRISM under out-domain conditions. Specifically, the combination of strata stratification with Lasso regularization (Str + L) achieved the best performance in most cohorts, with C-index values of 0.710, 0.712, 0.609, and 0.701 for GLCCM, AZCCM, RJCCM, and TJCCM, respectively. This finding underscores the importance of adaptive basis function learning in the CoxPH model for handling distribution shifts across different clinical centers.
@@ 3.5 Individualized PRISM-based risk interpretation
In contrast, among the listed complications strongly associated with MA-CE, case B exhibited two conditions, including microcirculatory disorder and ventricular aneurysm. Compared to MACE=0 samples, its differential EHR feature ranking showed less pronounced divergence than in case A, as reflected in the smaller Z-score deviation from zero. This observation may partly explain why PRISM adopted a more conservative approach in predicting MACE for case B.
@@ 4 Discussion
Integrating the cross-cohort SHAP value analyses derived from our PRISM model, we identified that among routinely obtained clinical assessments in standard hospital settings, hypertension and diabetes emerge as prominent potential biomarkers with high priority for predicting MACE. Notably, these comorbidities consistently ranked within the top-10 salient features across three of our independent centers. This finding underscores the persistent and critical role of traditional EHR variables alongside the advanced, unsupervised, image-based feature extraction facilitated by PRISM in survival analyses.
@@
5 Conclusion
PRISM demonstrates superior MACE survival prediction performance across four independent clinical cohorts, consistently outperforming classical and SOTA baselines under both internal and external validation. The framework’s unsupervised learning paradigm identifies three spatiotemporal imaging signatures, including lateral wall dyssynchrony, inferior wall hypersensitivity, and anterior diastolic elevated focus, that align with established coronary artery territories, validating its capacity to recapitulate pathophysiological principles without explicit anatomical supervision. Combined with prompt-guided attribution revealing hypertension, diabetes, and smoking as dominant EHR risk factors, PRISM establishes a practical, annotation-free approach for cardiovascular risk stratification that bridges high-dimensional imaging analytics with clinical assessments.
@@ 5 Conclusion
The study protocol was approved by the Institutional Review Board of Shanghai Renji Hospital (approval No. RJRE[2018]093). Collaborating centers, including Beijing Anzhen Hospital, Nanjing Drum Tower Hospital, and Wuhan Tongji Hospital, participated under this approved protocol. To protect patient privacy, informed consent was waived by the IRB above, and all data were anonymized prior to their use in model training, testing, and reader studies. This study was conducted in accordance with the Declaration of Helsinki and relevant local regulations.
Code availability
The source code used to generate the results presented in this study is available on GitHub at https://github.com/Hoyant-Su/PRISM. The repository includes scripts for data pre-processing, the three-stage training process, and evaluation of the model.
@@
H.S. designed the model and performed experiments. H.S. and S.R. developed and provided survival analysis code. H.S., S.R., and Y.G. processed the experimental data and documented control experiment details. J.X., X.C., T.Y., and L.W. provided cine imaging data, medical diagnoses, and electronic health records. S.Z., L.W., and X.W. supervised the research direction. H.S. drafted the manuscript. J.X., L.W., and X.W. contributed to manuscript editing and revision.
Corresponding author
Correspondence to Lian-Ming Wu and Xiaosong Wang.
Funding
This research received funding from the National Program for Support of Top-notch Young Professionals from the Organization Department of the Central Committee of the Communist Party of China, the National Natural Science Foundation of China (No. 82171884, No. 82471931), the Shanghai Municipal Commission of Science and Technology Medical Innovation Research Special Project (23Y11906900), and the Shanghai Municipal Health Commission Outstanding Youth Grant.